OpenStudy (anonymous):

|-x - 2y|. x is -2 and y is 3

OpenStudy (anonymous):

*evaluate.

OpenStudy (anonymous):

|-x -2y| = x-2y (without the negative sign in the answer)

OpenStudy (anonymous):

so...2 - 2(3)?

OpenStudy (anonymous):

Oh wait. i already tried that and the answer wasn't right.

OpenStudy (anonymous):

|-x-2y| = |-2 -2(-3)| = |-2+6| =

OpenStudy (anonymous):

ans is 4.

OpenStudy (anonymous):

@felavin Thank you. Would you mind telling me how you arrived at that answer, so I am able to learn? :)

OpenStudy (anonymous):

l-2+6l=l-4l = 4

OpenStudy (jhannybean):

@felavin work out problems instead of giving straight answers -.-

OpenStudy (anonymous):

Ohhh. Thank you, that helped. There's another question similar..|-x - 2y| when x is -5 and y is 5. I then got |-5 - 2(5)| and then got |-5 + 10|. My final answer was 5. Is that correct?

OpenStudy (anonymous):

ok.